# Domino tilings, nonintersecting lattice paths and subclasses of Koutschan–Krattenthaler–Schlosser determinants

### v1

@@ 2.2. Domino tilings
Let $M,N\in\mathbb{N}$ be natural numbers. We first introduce a domain $\mathfrak{A}_{\mathrm{D}}(M;N)$ as a prototype to facilitate our construction; this domain is built diagonal by diagonal. To start with, we work on a Cartesian grid and label each unit square by its left-top point. Now for each $k$ with $0\leq k\leq N$, we place $M+\lceil\frac{k}{2}\rceil$ unit squares, labeled by $(i,i-k)$ with $10\leq i\leq M+\lceil\frac{k}{2}\rceil-$ 1, into our domain $\mathfrak{A}_{\mathrm{D}}(M;N)$; these unit squares comprise the $k$-th diagonal of the domain. We also color even-indexed diagonals with white and odd-indexed diagonals with gray in the way of a chessboard. See Fig. 4 for an example.
@@
The domain involved in the counting function (${}_{k}(n)$ considered by Corteel, Huang and Krattenthaler is the Aztec-type domain of Type 1 marked by the partition $(k,k-1,\ldots,1,\underbrace{0,0,\ldots,0}_{n-k\text{ copies}})$.
@@
where #$\#_{T}(\mathfrak{D}_{i})$ counts the number of dominoes $\mathfrak{D}_{i}$ in the tiling $T$ for each $i=1,2,3,4$.
@@
(${}_{w_{1},w_{2},w_{3}}^{\lambda}(n)$ $:=\sum_{T\in\mydutchcal{D}_{\lambda}(n)}w_{\mathrm{D}}(T),$
^$\widehat{\cyD}_{w_{1},w_{2},w_{3}}^{\lambda}(n)$ $:=\sum_{\hat{T}\in\widehat{\mydutchcal{D}}_{\lambda}(n)}w_{\mathrm{D}}(\hat{T}).$
@@ 2.3. Bijections
#$\#_{T}(\mathfrak{D}_{1})=\#_{p}(\rightarrow),\qquad\#_{T}(\mathfrak{D}_{2})=\#_{p}(\uparrow),\qquad\#_{T}(\mathfrak{D}_{3})=\#_{p}(\nearrow).$
@@
#$\#_{\hat{T}}(\mathfrak{D}_{1})=\#_{\hat{p}}(\rightarrow),\qquad\#_{\hat{T}}(\mathfrak{D}_{2})=\#_{\hat{p}}(\uparrow),\qquad\#_{\hat{T}}(\mathfrak{D}_{3})=\#_{\hat{p}}(\nearrow).$
@@ 3. Delannoy and H-Delannoy determinants
Following the introduction of our domino tilings and nonintersecting lattice paths, now let us perform a short recitative to express the weighted enumerations (${}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ and ^$\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ for nonintersecting lattice paths as determinants involving weighted Delannoy and H-Delannoy numbers.
@@
(${}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ $=\underset{{0\leq i,j\leq n-1}}{\det}\big{(}D_{w_{1},w_{2},w_{3}}((s+1)i-j+r,j)\big{)},$ (3.5)
^$\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ $=\underset{{0\leq i,j\leq n-1}}{\det}\big{(}H_{w_{1},w_{2},w_{3}}((s+1)i-j+r,j)\big{)}.$ (3.6)
@@ 4. Subclasses of KKS determinants
Theorem 4.5 ($=$ Second Part in Theorems 1.1 and 1.2).
@@
Now on both sides, we replace $u$ with $e^{\frac{2\pi\mybbm{i}k}{m}}u^{\frac{1}{m}}$ for each $k$ with $10\leq k\leq m-$ 1 and sum over these $k$. Dividing the sum by $m$ and applying $\operatorname{RP}_{u}$, then in light of (4.1), the left-hand side becomes
@@
(${}_{1,l-1,1}^{(m-1,a)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left([u^{i}v^{j}]\frac{1-lv^{2}}{(1-v)(1-lv)}P_{1,l-1,1}^{(m-1,a)}\left(u,\frac{v}{(1-v)(1-lv)}\right)\right),$
@@ 5.1. Holonomic Ansatz
Now the takeaway of the holonomic Ansatz principle is that by defining the $(n-1,j)$-th normalized cofactors of $A_{n}$:
@@
respectively, where $S$ represents the forward shift operator defined by $S_{l}\circ f(l):=f(l+1)$ for functions $f$ in the variable $l$. These annihilators occupy 819 944819944819\,944819 944 Mathematica bytes in total.
@@
and occupies 200 760200760200\,760200 760 Mathematica bytes.
@@ 5.2.1. Relation (H3)
The above computation for modular reduction was executed in parallel on an Amazon cluster (AWS EC2, instance type: m5.24xlarge [^1] See https://instances.vantage.sh/aws/ec2/m5.24xlarge for details., 96 CPUs, 384 GB Memory [^2] We report that a Memory overflow problem was always encountered due to the scale of our computation at our initial attempts so eventually we had to launch only 48 kernels for interpolations and 72 kernels for rational reconstruction.). In particular, it took around 1.5 Days to generate interpolating points and another 2.5 Days for the rational reconstruction of all $21+691=712$ parameters. The outputting telescoper part occupies 6 102 52861025286\,102\,5286 102 528 Mathematica bytes and the delta part occupies 3 182 669 57631826695763\,182\,669\,5763 182 669 576 bytes. The scale of them, which is extraordinarily huge, gives the second (and the most critical) reason why an automatic application of Koutschan’s FindCreativeTelescoping command is impractical. For the data of other calculations, we refer the reader to Table 2.
@@ 5.2.2. Relation (H2)
where $10\leq i<n-$ 1. As before, we apply creative telescoping under modular reduction for the two summands $a_{i,j}^{(1)}c_{n,j}$ and $a_{i,j}^{(2)}c_{n,j}$, respectively, and we will still use the former as an instance.
@@
In practice, we choose $K=160$ different $i$-values, $L=528$ different $n$-values, and $M=40$ different large primes of size around $2^{32}$. Finally, the rational function $P(i,n)$ can be reconstructed by these interpolations. As soon as we obtain all unknown parameters in the telescoper and delta parts, we arrive at a conjectural annihilating operator (5.9) for $a_{i,j}^{(1)}c_{n,j}$, which can be certified by the OreReduce command of the HolonomicFunctions package. The whole computation was executed parallelly on our Amazon cluster, and it took around 1.5 Days to generate interpolating points and another 1 Day for the rational reconstruction of all $21+148=169$ parameters. The outputting telescoper part occupies 27 393 8082739380827\,393\,80827 393 808 Mathematica bytes and the delta part occupies 1 777 395 42417773954241\,777\,395\,4241 777 395 424 bytes. We remark that by the theory of creative telescoping, the telescoper part annihilates $\Sigma_{i,n}^{(1)}$. In the same way, we may also obtain an annihilator for $\Sigma_{i,n}^{(2)}$ supported on $\operatorname{Supp}_{\mathrm{H}_{2}}^{\dagger}$. It is easy to check that they differ by a constant scalar, and hence either of them annihilates $\Sigma_{i,n}$. We denote such an annihilator for $\Sigma_{i,n}$ by $\mathsf{annH2a}$.
@@
$\operatorname{Ind}_{\mathrm{H}_{2}}^{\dagger}$ :=$:=\big{\{}(a,b):S_{i}^{a}S_{n}^{b}\in\operatorname{Supp}_{\mathrm{H}_{2}}^{\dagger}\backslash\{S_{i}^{5}\}\big{\}},$
$\operatorname{Ind}_{\mathrm{H}_{2}}^{\ddagger}$ :=$:=\big{\{}(a,b):S_{i}^{a}S_{n}^{b}\in\operatorname{Supp}_{\mathrm{H}_{2}}^{\ddagger}\backslash\{S_{i}^{6}\}\big{\}}.$
@@
Finally, we notice that $a_{i_{0},j}^{(2)}=\binom{4j-i_{0}}{4j+2}$ equals $\binom{-i_{0}}{2}$ when $j=0$ and 00 whenever $j\geq 1$. Hence,
@@ 6. Epilogue
Before bringing down the curtain, we revisit the question of Koutschan, Krattenthaler and Schlosser, and think about other connections between KKS determinants and our weighted enumerations of domino tilings, (${}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$, or equivalently, nonintersecting lattice paths, (${}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$. Based on a brief computer search for relations concerning (${}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ and (${}_{a,b,c,d}^{(m,l)}(n)$, we find that apart from (1.10), it seems necessary to require one of $w_{2}$ and $w_{3}$ to be zero when we set $s$ and $r$ generic. Note that the case where $w_{1}=0$ is not at all interesting because there always exists an east step in the Delannoy path from $(-n,n)$ to $(r-n,n)$ for $r\geq 1$ so that the weighted contribution from the associated system of nonintersecting Delannoy paths becomes zero. In view of Remark 2.2, for the moment we only look at (${}_{1,1,0}^{(s,r)}(n)$ and (${}_{1,0,1}^{(s,r)}(n)$.
@@
(${}_{1,0,1}^{(s,r)}(n)={}_{r+2\varrho,b,r+2\varrho,r}^{(s+1,1)}(n)=\tfrac{1}{2}{}_{r,b,r,r}^{(s+1,1)}(n).$ (6.2)
@@
For the evaluation of (${}_{0,0,c,c}^{(1,s+2)}(n)$, we directly make use of a result of Koutschan, Krattenthaler and Schlosser [16, p. 6, Theorem 3] to get
(${}_{0,0,c,c}^{(1,s+2)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left((s+2)^{j}\binom{i+j+c}{i}+\binom{i-j+c}{i}\right)=2(s+1)^{\binom{n}{2}}.$
For the evaluation of (${}_{1,1,0}^{(s,r)}(n)$, we first recall from (3.5) that
(${}_{1,1,0}^{(s,r)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\big{(}D_{1,1,0}((s+1)i-j+r,j)\big{)}.$
@@
which follows from either the double series expansion of (3.3) or the fact that $D_{1,1,0}((s+1)i-j+r,j)$ indeed counts the number of paths from $(0,0)$ to $((s+1)i-j+r,j)$ by using only east and north steps. Therefore,
(${}_{1,1,0}^{(s,r)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left(\binom{(s+1)i+r}{j}\right).$
@@
(${}_{a,b,a,d}^{(m,1)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left(\binom{mi+j+a}{mi+a}+\binom{mi-j+d}{mi+a}\right),$
@@
(${}_{a+2\varrho,b,a+2\varrho,a}^{(m,1)}(n)=\tfrac{1}{2}{}_{a,b,a,a}^{(m,1)}(n).$ (6.3)
@@
(${}_{a,b,a,a}^{(m,1)}(i,j)-\varepsilon(i,j)=\begin{cases}\tfrac{1}{2}{}_{a,b,a,a}^{(m,1)}(i,j),&\text{if $j=0$},\\[6.0pt] {}_{a,b,a,a}^{(m,1)}(i,j),&\text{if $j\geq 1$}.\end{cases}$
@@
(~$\big{(}\widetilde{\cyb}_{a+2\varrho,b,a+2\varrho,a}^{(m,1)}(i,j)\big{)}_{i,j\geq 0}:=\mydutchcal{B}_{a+2\varrho,b,a+2\varrho,a}^{(m,1)}\cdot\big{(}(-1)^{j-i}\tbinom{2\varrho}{j-i}\big{)}_{i,j\geq 0}.$
@@
^$\widehat{\cyL}_{1,1,0}^{(s,r)}(n)=\tfrac{1}{2}{}_{0,0,c,c}^{(1,s+2)}(n)=(s+1)^{\binom{n}{2}}.$ (6.8)
@@
^$\widehat{\cyL}_{1,0,1}^{(s,r)}(n)={}_{r+2\varrho-1,b,r+2\varrho-1,r-1}^{(s+1,1)}(n)=\tfrac{1}{2}{}_{r-1,b,r-1,r-1}^{(s+1,1)}(n).$ (6.9)
@@
It is clear from (3.3) and (3) that $H_{1,1,0}(i,j)=D_{1,1,0}(i,j)$; this can also be seen combinatorially from the facts that for each H-Delannoy path from $(-j,j)$ to $(s(n-j)+r-j,n+1)$, the position $(s(n-j)+r-j-1,n+1)$ is forbidden by definition and that all northeast steps are also not allowed since the weight $w_{3}=0$, so that this H-Delannoy path must pass through $(s(n-j)+r-j,n)$ and hence becomes a Delannoy path ending at $(s(n-j)+r-j,n)$ with the weights for the three types of steps being $(w_{1},w_{2},w_{3})=(1,1,0)$ as assumed. Therefore, (6.8) comes directly from (6.1).
@@
(${}_{r+2\varrho-1,b,r+2\varrho-1,r-1}^{(s+1,1)}(n)=\tfrac{1}{2}{}_{r-1,b,r-1,r-1}^{(s+1,1)}(n)$

### v2

@@ 2.2. Domino tilings
Let $M,N\in\mathbb{N}$ be natural numbers. We first introduce a domain $\mathfrak{A}_{\mathrm{D}}(M;N)$ as a prototype to facilitate our construction; this domain is built diagonal by diagonal. To start with, we work on a Cartesian grid and label each unit square by its left-top point. Now for each $k$ with $0\leq k\leq N$, we place $M+\lceil\frac{k}{2}\rceil$ unit squares, labeled by (i,ik)(i,i-k) with $0\leq i\leq M+\lceil\frac{k}{2}\rceil-1$, into our domain $\mathfrak{A}_{\mathrm{D}}(M;N)$; these unit squares comprise the $k$-th diagonal of the domain. We also color even-indexed diagonals with white and odd-indexed diagonals with gray in the way of a chessboard. See Fig. 4 for an example.
@@
[…]
@@
where $_{T}(\mathfrak{D}_{i})$ counts the number of dominoes $\mathfrak{D}_{i}$ in the tiling $T$ for each $i=1,2,3,4$.
@@
[…]
@@ 2.3. Bijections
$\#_{T}(\mathfrak{D}_{1})=\#_{p}(\rightarrow),\qquad\#_{T}(\mathfrak{D}_{2})=\#_{p}(\uparrow),\qquad\#_{T}(\mathfrak{D}_{3})=\#_{p}(\nearrow).$
@@
[…]
@@ 3. Delannoy and H-Delannoy determinants
Following the introduction of our domino tilings and nonintersecting lattice paths, now let us perform a short recitative to express the weighted enumerations $_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ and $\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ for nonintersecting lattice paths as determinants involving weighted Delannoy and H-Delannoy numbers.
@@
${}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ $=\underset{{0\leq i,j\leq n-1}}{\det}\big{(}D_{w_{1},w_{2},w_{3}}((s+1)i-j+r,j)\big{)},$ (3.5)
$\widehat{\cyL}_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ $=\underset{{0\leq i,j\leq n-1}}{\det}\big{(}H_{w_{1},w_{2},w_{3}}((s+1)i-j+r,j)\big{)}.$ (3.6)
@@ 4. Subclasses of KKS determinants
Theorem 4.5 (== Second Part in Theorems 1.1 and 1.2).
@@
Now on both sides, we replace $u$ with $e^{\frac{2\pi\mybbm{i}k}{m}}u^{\frac{1}{m}}$ for each $k$ with $0\leq k\leq m-1$ and sum over these $k$. Dividing the sum by $m$ and applying $\operatorname{RP}_{u}$, then in light of (4.1), the left-hand side becomes
@@
${}_{1,l-1,1}^{(m-1,a)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left([u^{i}v^{j}]\frac{1-lv^{2}}{(1-v)(1-lv)}P_{1,l-1,1}^{(m-1,a)}\left(u,\frac{v}{(1-v)(1-lv)}\right)\right),$
@@ 5.1. Holonomic Ansatz
Now the takeaway of the holonomic Ansatz principle is that by defining the (n1,j)(n-1,j)-th normalized cofactors of $A_{n}$:
@@
respectively, where $S$ represents the forward shift operator defined by $S_{l}\circ f(l):=f(l+1)$ for functions $f$ in the variable $l$. These annihilators occupy 819 944819\,944 Mathematica bytes in total.
@@
and occupies 200 760200\,760 Mathematica bytes.
@@ 5.2.1. Relation (H3)
The above computation for modular reduction was executed in parallel on an Amazon cluster (AWS EC2, instance type: m5.24xlarge [^1] See https://instances.vantage.sh/aws/ec2/m5.24xlarge for details., 96 CPUs, 384 GB Memory [^2] We report that a Memory overflow problem was always encountered due to the scale of our computation at our initial attempts so eventually we had to launch only 48 kernels for interpolations and 72 kernels for rational reconstruction.). In particular, it took around 1.5 Days to generate interpolating points and another 2.5 Days for the rational reconstruction of all $21+691=712$ parameters. The outputting telescoper part occupies 6 102 5286\,102\,528 Mathematica bytes and the delta part occupies 3 182 669 5763\,182\,669\,576 bytes. The scale of them, which is extraordinarily huge, gives the second (and the most critical) reason why an automatic application of Koutschan’s FindCreativeTelescoping command is impractical. For the data of other calculations, we refer the reader to Table 2.
@@ 5.2.2. Relation (H2)
where $0\leq i<n-1$. As before, we apply creative telescoping under modular reduction for the two summands $a_{i,j}^{(1)}c_{n,j}$ and $a_{i,j}^{(2)}c_{n,j}$, respectively, and we will still use the former as an instance.
@@
In practice, we choose $K=160$ different $i$-values, $L=528$ different $n$-values, and $M=40$ different large primes of size around $2^{32}$. Finally, the rational function $P(i,n)$ can be reconstructed by these interpolations. As soon as we obtain all unknown parameters in the telescoper and delta parts, we arrive at a conjectural annihilating operator (5.9) for $a_{i,j}^{(1)}c_{n,j}$, which can be certified by the OreReduce command of the HolonomicFunctions package. The whole computation was executed parallelly on our Amazon cluster, and it took around 1.5 Days to generate interpolating points and another 1 Day for the rational reconstruction of all $21+148=169$ parameters. The outputting telescoper part occupies 27 393 80827\,393\,808 Mathematica bytes and the delta part occupies 1 777 395 4241\,777\,395\,424 bytes. We remark that by the theory of creative telescoping, the telescoper part annihilates $\Sigma_{i,n}^{(1)}$. In the same way, we may also obtain an annihilator for $\Sigma_{i,n}^{(2)}$ supported on $\operatorname{Supp}_{\mathrm{H}_{2}}^{\dagger}$. It is easy to check that they differ by a constant scalar, and hence either of them annihilates $\Sigma_{i,n}$. We denote such an annihilator for $\Sigma_{i,n}$ by $\mathsf{annH2a}$.
@@
[…]
@@
Finally, we notice that $a_{i_{0},j}^{(2)}=\binom{4j-i_{0}}{4j+2}$ equals $\binom{-i_{0}}{2}$ when $j=0$ and 0 whenever $j\geq 1$. Hence,
@@ 6. Epilogue
Before bringing down the curtain, we revisit the question of Koutschan, Krattenthaler and Schlosser, and think about other connections between KKS determinants and our weighted enumerations of domino tilings, $_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$, or equivalently, nonintersecting lattice paths, $_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$. Based on a brief computer search for relations concerning $_{w_{1},w_{2},w_{3}}^{(s,r)}(n)$ and $_{a,b,c,d}^{(m,l)}(n)$, we find that apart from (1.10), it seems necessary to require one of $w_{2}$ and $w_{3}$ to be zero when we set $s$ and $r$ generic. Note that the case where $w_{1}=0$ is not at all interesting because there always exists an east step in the Delannoy path from (n,n)(-n,n) to (rn,n)(r-n,n) for $r\geq 1$ so that the weighted contribution from the associated system of nonintersecting Delannoy paths becomes zero. In view of Remark 2.2, for the moment we only look at $_{1,1,0}^{(s,r)}(n)$ and $_{1,0,1}^{(s,r)}(n)$.
@@
${}_{1,0,1}^{(s,r)}(n)={}_{r+2\varrho,b,r+2\varrho,r}^{(s+1,1)}(n)=\tfrac{1}{2}{}_{r,b,r,r}^{(s+1,1)}(n).$ (6.2)
@@
For the evaluation of $_{0,0,c,c}^{(1,s+2)}(n)$, we directly make use of a result of Koutschan, Krattenthaler and Schlosser [16, p. 6, Theorem 3] to get
${}_{0,0,c,c}^{(1,s+2)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left((s+2)^{j}\binom{i+j+c}{i}+\binom{i-j+c}{i}\right)=2(s+1)^{\binom{n}{2}}$.
For the evaluation of $_{1,1,0}^{(s,r)}(n)$, we first recall from (3.5) that
${}_{1,1,0}^{(s,r)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\big{(}D_{1,1,0}((s+1)i-j+r,j)\big{)}.$
@@
which follows from either the double series expansion of (3.3) or the fact that $D_{1,1,0}((s+1)i-j+r,j)$ indeed counts the number of paths from $(0,0)$ to $($s+1)ij+r,j)$($s+1)i-j+r,j) by using only east and north steps. Therefore,
${}_{1,1,0}^{(s,r)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left(\binom{(s+1)i+r}{j}\right)$.
@@
${}_{a,b,a,d}^{(m,1)}(n)=\underset{{0\leq i,j\leq n-1}}{\det}\left(\binom{mi+j+a}{mi+a}+\binom{mi-j+d}{mi+a}\right)$,
@@
${}_{a+2\varrho,b,a+2\varrho,a}^{(m,1)}(n)=\tfrac{1}{2}{}_{a,b,a,a}^{(m,1)}(n).$ (6.3)
@@
${}_{a,b,a,a}^{(m,1)}(i,j)-\varepsilon(i,j)=\begin{cases}\tfrac{1}{2}{}_{a,b,a,a}^{(m,1)}(i,j),&\text{if $j=0$},\\[6.0pt] {}_{a,b,a,a}^{(m,1)}(i,j),&\text{if $j\geq 1$}.\end{cases}$
@@
$\big{(}\widetilde{\cyb}_{a+2\varrho,b,a+2\varrho,a}^{(m,1)}(i,j)\big{)}_{i,j\geq 0}:=\mydutchcal{B}_{a+2\varrho,b,a+2\varrho,a}^{(m,1)}\cdot\big{(}(-1)^{j-i}\tbinom{2\varrho}{j-i}\big{)}_{i,j\geq 0}.$
@@
$\widehat{\cyL}_{1,1,0}^{(s,r)}(n)=\tfrac{1}{2}{}_{0,0,c,c}^{(1,s+2)}(n)=(s+1)^{\binom{n}{2}}$. (6.8)
@@
$\widehat{\cyL}_{1,0,1}^{(s,r)}(n)={}_{r+2\varrho-1,b,r+2\varrho-1,r-1}^{(s+1,1)}(n)=\tfrac{1}{2}{}_{r-1,b,r-1,r-1}^{(s+1,1)}(n).$ (6.9)
@@
It is clear from (3.3) and (3) that $H_{1,1,0}(i,j)=D_{1,1,0}(i,j)$; this can also be seen combinatorially from the facts that for each H-Delannoy path from (j,j)(-j,j) to (s(nj)+rj,n+1)(s(n-j)+r-j,n+1), the position (s(nj)+rj1,n+1)(s(n-j)+r-j-1,n+1) is forbidden by definition and that all northeast steps are also not allowed since the weight $w_{3}=0$, so that this H-Delannoy path must pass through (s(nj)+rj,n)(s(n-j)+r-j,n) and hence becomes a Delannoy path ending at (s(nj)+rj,n)(s(n-j)+r-j,n) with the weights for the three types of steps being $(w_{1},w_{2},w_{3})=(1,1,0)$ as assumed. Therefore, (6.8) comes directly from (6.1).
@@
${}_{r+2\varrho-1,b,r+2\varrho-1,r-1}^{(s+1,1)}(n)=\tfrac{1}{2}{}_{r-1,b,r-1,r-1}^{(s+1,1)}(n)$